The length of a rectangle is 15 cm more than its width. A second rectangle whose perimeter is 72 cm is 5 cm wider but 2 cm shorter than the first rectangle. What are the dimensions of each rectangle?
Answer by mananth(16946)

FIRST RECTANGLE
let x be the width
length = x+15
...
SECOND RECTANGLE
5 cm wider = x+5
2cm shorter = x+15-2=x+13
..
2*(l+w)=perimeter
2*((x+5)+(x+13))=72
2*(x+5+x+13)=72
2*(2x+18)=72
divide by 2
2x+18= 72/2
2x+18=36
-18 -18
2x+18=18=36-18
2x=18
/2
x=18/2
x=9 the width
length = x+15 = 9+15=24
FIRST RECTANGLE
WIDTH 9 CM
LENGTH = 24
SECOND RECTANGLE
WIDTH = x+5 = 9+5=14
LENGTH = x+13= 9+13=22
...
CHECK
2*14+2*22=72
...
m.ananth@hotmail.ca